two chunks of dry ice rest on a smooth table top. one has a mass of 1 kg, the other is 2 kg. A 4.0 N force is applied to each for 4 s.

a. What is the acceleration of each chunk?
b. Compare the velocities of the two chunks at the end of 4s.
c. Compare the displacement of the two chunks at the end of 4s.
Answer #1

For ice of mass 1 kg

F = m*a

a = F/m = 4/1 = 4 m/sec^2

V = u + a*t

V = 0 + 4*4 = 16 m/sec

d = u*t + 0.5*a*t^2

d = 0*4 + 0.5*4*4^2 = 32 m

For ice of mass 2 kg

F = m*a

a = F/m = 4/2 = 2 m/sec^2

V = u + a*t

V = 0 + 2*4 = 8 m/sec

d = u*t + 0.5*a*t^2

d = 0*4 + 0.5*2*4^2 = 16 m